# Monte Carlo Tree Diffusion with Multiple Experts for Protein Design

### v1

@@ 6 CONCLUSION
In conclusion, we introduced MCTD-ME, which combines diffusion-based generative planning with MCTS and multi-expert guidance. The method is novel in reframing diffusion decoding as tree search for parallel multi-token exploration and in integrating multiple expert evaluators with a pLDDT-guided masking schedule to improve biophysical fidelity. Our selection policy ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) efficiently arbitrates among experts using cached uncertainty statistics, enabling stable and low-overhead rollouts. Empirically, MCTD-ME outperforms state-of-the-art baselines on inverse folding, scales effectively with additional computation, and provides a general template for diffusion-plus-search workflows in structured generative tasks.
@@ Appendix A Appendix
A.1 Algorithm details
@@
1:
[…]
4:$K$: number of children per expansion
5:$E$: number of expert models
6:$R$: rollouts per expert
7:$T$: total simulations
@@
11:for $i=1$ to $T$ do
12:  $node\leftarrow root$
13:  $\triangleright$ /* Selection via $\text{P}\mathcal{H}\text{-UCT}\text{-ME}*$/
14:  while $node.children\neq\emptyset$ do
15:   $node\leftarrow\text{P}\mathcal{H}\text{-UCT}\text{-ME}(node.children)$ // uses $Q$, UCB, and cached $(\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}})$
16:  $\triangleright$ /* Expansion via pLDDT Masking and Multi-Expert Rollouts */
17:  $\mathcal{M}\leftarrow\textsc{GetMaskSet}(node.sequence)$ // Based on pLDDT
18:  $\mathcal{Y}\leftarrow\emptyset$ // candidate children
19:  for $e\in\mathcal{E}$ do // each generator expert
20:   for $r=1$ to $R$ do // rollouts
21:     $y_{0}\leftarrow\textsc{MaskedDiffusion}(node.sequence,\mathcal{M},e)$
22:     if $y_{0}\notin cache$ then
23:      $score\leftarrow\textsc{EvalComposite}(y_{0};\mathcal{C})$ // weighted sum of critics
24:      $cache[y_{0}]\leftarrow score$
25:     $\mathcal{Y}\leftarrow\mathcal{Y}\cup\{y_{0}\}$
26:  $\mathcal{Y}_{\text{top}}\leftarrow\textsc{TopK}(\mathcal{Y},K;\,cache)$ // sort by composite reward
27:  for $y^{\prime}\in\mathcal{Y}_{\text{top}}$ do
28:   $(\mathcal{U}_{\text{ent}}(y^{\prime}),\mathcal{U}_{\text{div}}(y^{\prime}))\leftarrow\textsc{ComputeBonuses}(y^{\prime},\,node.sequence,\,\mathcal{M})$ // ensemble surprisal & novelty
29:   $\textsc{AddChild}(node,y^{\prime},\,\mathcal{U}_{\text{ent}}(y^{\prime}),\,\mathcal{U}_{\text{div}}(y^{\prime}))$ // create new node with full seq
30:  $\triangleright$ /* Evaluation of New Children (full-sequence) */
31:  for $y^{\prime}\in\mathcal{Y}_{\text{top}}$ do
32:   $v\leftarrow cache[y^{\prime}]$ // may hit cache; same composite reward as above
33:   $\triangleright$ /* Backpropagation */
34:   $\textsc{Backpropagate}(y^{\prime},\,v)$ // supports max or sum backup
35:return Top-$k$ sequences by $cache[\cdot]$
A.2 Dataset Details
We evaluate MCTD-ME on two protein sequence benchmarks at the chain level. (1) CAMEO2022 comprises 183 targets released in 2022; we use the reference backbones and evaluate inverse folding on the corresponding chains. (2) PDB_date is a date-split benchmark of 7,855 proteins constructed from the Protein Data Bank (PDB); we hold out targets by deposition date and evaluate inverse folding on their chains. In our experiments we use a date-split subset of 449 chains from PDB_date for compute efficiency; unless otherwise noted, length statistics below are for the full benchmark. For both sets, sequence lengths are computed after simple canonicalization (upper-casing and removal of non-canonical/unknown residues).

### v2

@@ 6 Conclusion
We presented MCTD-ME, a model-agnostic planner that reframes reverse diffusion as Monte Carlo tree search, enabling parallel multi-token revisions and principled exploration guided by multiple proposal experts. A pLDDT-driven masking schedule focuses denoising on structurally uncertain regions, improving biophysical fidelity, while our $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection reuses cached uncertainty to efficiently arbitrate among experts without recomputing entropies. Across folding, inverse folding, and motif scaffolding, MCTD-ME consistently outperforms single-expert and one-shot baselines—raising AAR and scTM and lowering RMSD—with gains that scale with test-time compute and are robust to the choice of experts. Although the method currently relies on surrogate structure predictors and incurs some search overhead as a limitation, it provides a unified, plug-and-play test-time mechanism that reliably improves design quality. Future work will integrate learned critics and apply further systems optimizations to enhance both accuracy and speed.
@@ Appendix A Appendix
A.1 Algorithm details.
Lead optimiaztion
In this algorithm, we use MCTD-ME to tackle the lead optimization challenge. Beginning with an initial lead—a fully denoised baseline sequence—we execute MCTS with the following procedure: (i) selection uses a $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$–style rule that combines value (Q), exploration, and cached uncertainty bonuses (ensemble surprisal ($\mathcal{U}_{\text{ent}}$) and novelty ($\mathcal{U}_{\text{div}}$)); (ii) expansion proposes edits only on pLDDT-masked uncertain sites and performs multi-expert rollouts (E generators), producing candidates scored once via a composite critic (cached to avoid re-evaluation); (iii) we keep the Top-K children by composite reward, attach their uncertainty bonuses, and (iv) backpropagate the candidate’s value to update ancestors (max/sum backup). This yields a planner that routes among experts, focuses edits where structure is uncertain, reuses evaluations via caching, and optimizes a task-aware composite objective.
@@
[…]
De novo generation
We start from an all-mask root. At each iteration, Selection ($\text{P}\mathcal{H}\text{-UCT}\text{-ME}$) walks to a leaf. Expansion proposes partial fills only at low-pLDDT sites (high-pLDDT positions are frozen). Then Evaluation performs short rollouts from each new child (e.g., a few masked-diffusion steps / fast fold+critic passes) to estimate a terminal composite value; results are cached. Finally, we Backpropagate those values. This keeps the planner focused on uncertain regions while using cheap rollouts to decide which partials to keep; see Algorithm 2 for pseudocode.
Algorithm 2 De Novo MCTD with Multiple Experts (MCTD-ME -DeNovo)
[…]
A.2 Dataset details.
We evaluate MCTD-ME on two protein sequence benchmarks at the chain level. As illustrated in Table 4 (1) CAMEO2022 comprises 183 targets released in 2022; we use the reference backbones and evaluate inverse folding on the corresponding chains. (2) PDB_date is a date-split benchmark of 7,855 proteins constructed from the Protein Data Bank (PDB); we hold out targets by deposition date and evaluate inverse folding on their chains. In our experiments we use a date-split subset of 449 chains from PDB_date for compute efficiency; unless otherwise noted, length statistics below are for the full benchmark. For both sets, sequence lengths are computed after simple canonicalization (upper-casing and removal of non-canonical/unknown residues).